### v1

# GRAFT: Gradient-Aware Fast MaxVol Technique for Dynamic Data Sampling

Abstract
Training modern neural networks on large datasets is computationally and environmentally costly. We introduce GRAFT, a scalable in-training subset selection method that (i) extracts a low-rank feature representation for each per-iteration batch, (ii) applies a Fast MaxVol sampler to pick a small, diverse subset that spans the batch’s dominant subspace, and (iii) dynamically adjusts the subset size using a gradient-approximation criterion. By operating in low-rank subspaces and training on carefully chosen examples instead of full batches, GRAFT preserves the training trajectory while reducing wall-clock time, energy, and $\mathrm{CO}_{2}$ emissions. Across multiple benchmarks, it matches or exceeds recent selection baselines in accuracy and efficiency, while providing a favorable accuracy efficiency emissions trade-off.
1 Introduction
Deep neural networks are powerful but costly to train, requiring substantial computational resources and energy, raising scalability and environmental concerns. A key driver of this cost is the large volume of training data, which increases optimization steps, memory use, and training time. Data subset selection offers a solution by using a smaller, well-chosen subset to maintain training efficiency while reducing redundancy. This can speed up training, cut energy use, and aid resource-constrained settings. However, many existing methods have drawbacks, such as costly preprocessing, reliance on proxy models, architecture restrictions, or repeated metric evaluations, limiting their flexibility.
In this work, we introduce GRAFT (Gradient-Aware Fast MaxVol Technique), a scalable and lightweight framework designed to integrate subset selection directly into the training loop. GRAFT operates in two main stages - first, it extracts compact feature embeddings from per-iteration batch using a low-rank projection. This step ensures that the high-dimensional representation of the data is mapped onto a lower-dimensional subspace, capturing the most salient features with minimal redundancy. Next, GRAFT performs a fast MaxVol sampling to select a subset of examples that best span the dominant subspace of the projected features. Unlike heuristic-based selection strategies, the MaxVol criterion guarantees that the chosen subset retains maximal expressiveness in terms of subspace coverage. To maintain alignment with the training dynamics, GRAFT introduces an adaptive sampling mechanism that dynamically adjusts the number of selected samples per-iteration batch. This adjustment is guided by measuring the alignment between the gradient computed over the per-iteration subset and the gradient of the entire per-iteration batch. Specifically, GRAFT evaluates the subspace representation of the selected subset to ensure it remains representative of the full batch’s learning dynamics. If the alignment deviates, GRAFT automatically increases the subset size to better capture the gradient direction. Conversely, when alignment is strong, GRAFT reduces the subset size to optimize efficiency. This allows GRAFT to preserve training trajectory fidelity, ensuring that critical gradient information is not lost during optimization, even as subset compositions evolve throughout training. Furthermore, GRAFT introduces a fundamentally different mechanism compared to existing subset selection strategies such as GradMatch. GradMatch focuses on gradient matching, where the primary goal is to select a subset $\mathcal{S}$ such that the gradient computed over $\mathcal{S}$ closely approximates the gradient of the entire dataset. This is achieved through a greedy orthogonal matching pursuit (OMP) strategy that incrementally selects samples to minimize the residual gradient error. While effective, gradient matching methods require explicit comparisons of gradient vectors, which can be computationally expensive and sensitive to noise.
In contrast, GRAFT shifts the focus from direct gradient matching to per-iteration subspace approximation combined with gradient alignment. Importantly, GRAFT ensures that this subspace representation remains gradient-aligned with the full per-iteration batch by dynamically adjusting the subset size based on the measured alignment. This mechanism guarantees that the chosen samples not only capture the most informative directions in the data but also maintain the critical gradient signals necessary for effective optimization. As a result, GRAFT achieves improved efficiency by reducing reliance on full-gradient computations while preserving the quality of the training trajectory, leading to more robust and stable convergence.
Figure 1: Overview of the GRAFT sampling framework. The method consists of two main stages: (1) Feature Extraction and Sample Selection, where input per-iteration batch $X$ is transformed into a low-dimensional feature matrix $Z=f(X)$, followed by MaxVol-based selection of representative rows (samples); and (2) Gradient Alignment and Rank Selection, where the selected subset is evaluated based on how well its gradient aligns with the full per-iteration batch gradient. The selection rank $R$ is dynamically adjusted using a projection error criterion. The final output is a subset of samples indexed by $I=\{i_{1},\ldots,i_{R}\}$ used for gradient updates.
This design enables GRAFT to operate efficiently within small localized data, using the local structure to guide training. By selecting representative samples in each iteration, GRAFT reduces the computational effort in subset selection while adapting to the evolving optimization landscape throughout the training. Our experimental results demonstrate that GRAFT achieves strong performance across multiple datasets and architectures, reducing both training time and energy usage with minimal accuracy degradation. These findings position GRAFT as a practical tool for sustainable and efficient deep learning on a scale.
2 Related Works
Subset selection methods can be broadly categorized into pre-selection and integrated selection approaches. Pre-selection methods identify representative subsets before training. A notable example is EL2N (Early-Loss-to-Norm) scoring, defined as $\text{EL2N}(x_{i})=\frac{1}{t}\sum_{k=1}^{t}|\nabla L(\theta_{k};x_{i})|^{2}$, which leverages early-phase gradient norms to approximate sample informativeness [1]. Samples with higher early gradient norms tend to have a stronger impact on parameter updates and loss convergence. Forget Score [2] adopts a different strategy, measuring the misclassification frequency of samples as $F(x_{i})=\sum_{k=1}^{t}\mathbb{1}[y_{i}\neq\hat{y}_{i}^{(k)}]$, which favors harder-to-learn examples. However, this method is highly sensitive to noise and class imbalance, potentially overemphasizing mislabeled or ambiguous samples. Selection via Proxy (SVP) [3] introduces a lightweight proxy model $f_{\phi}$ trained independently to estimate sample informativeness, formulated as $\text{SVP}(x_{i})=L(f_{\phi}(x_{i}),y_{i})$. Although efficient, its effectiveness is tightly coupled to the fidelity of the proxy model, which may not always generalize well to the main model’s learning dynamics. DRoP [4] conceptualizes subset selection as a distributionally robust optimization (DRO) problem, expressed as $\min_{\theta}\max_{q\in\mathcal{Q}}\sum_{i=1}^{n}q_{i}L(f_{\theta}(x_{i}),y_{i})$. This adversarial perspective enhances robustness against noise, but it introduces significant computational overhead due to adversarial evaluations during each optimization step. One inherent limitation of pre-selection methods is their inability to adapt to evolving training dynamics, often resulting in underfitting or misalignment with the model’s learning trajectory. In contrast, integrated selection methods address this shortcoming by adjusting subsets throughout training, aligning selection with gradient-aligned subspace approximations that reflect the current optimization landscape. This adaptability is especially critical in non-convex optimization settings where data importance fluctuates across iterations. Prominent examples include GradMatch [5], which minimizes the gradient discrepancy $|\nabla_{\theta}\mathcal{L}(S)-\nabla_{\theta}\mathcal{L}(\mathcal{X})|_{2}^{2}$. Despite its theoretical alignment with full-batch gradients, GradMatch suffers from substantial computational overhead during optimization. GLISTER [6] and CRAIG [7] frame the subset selection problem as a submodular optimization task, which allows efficient greedy selection but lacks adaptive flexibility to dynamic gradient changes. GRAFT builds upon these limitations by improving scalability and optimization fidelity. It samples from a low-rank subspace, maximizing the spanned volume through Fast MaxVol-based selection, and dynamically adjusts subset sizes to preserve representational capacity throughout training. Empirical evaluations using the CORDS framework demonstrate that methods like GradMatch and GLISTER consistently encounter out-of-memory errors with large batch sizes, highlighting their scalability constraints. Other relevant techniques include BADGE [8], which combines uncertainty with gradient embedding clustering to target informative samples. Moderately hard core-sets [9] aim to select samples of intermediate difficulty, striking a balance between learnability and informativeness. Dataset distillation [10] attempts to synthesize compact datasets but struggles to maintain representativeness at moderate subset sizes. Hybrid strategies like SelMatch [11] refine real data selections through gradient matching, conceptually similar to GRAFT, although they require additional meta-optimization layers. SubSelNet [12], designed specifically for NAS, learns generalized selection policies. However, its reliance on prior task-architecture distributions constrains its generalizability to out-of-distribution scenarios.
Method Batch-wise Gradient-based Complexity Scalability Bottleneck Dynamic Rank Memory Usage
DRoP $\mathcal{O}(n^{2}d)$ Quadratic in $n$ (intractable for $n>10^{4}$) High
GradMatch $\mathcal{O}(nkd)$ Linear in $n$ (full gradient comparisons) High
CRAIG $\mathcal{O}(nkd)$ Linear in $n$ (submodular optimization) Moderate
GLISTER $\mathcal{O}(nkd)$ Linear in $n$ (bi-level optimization) Moderate
SVP $\mathcal{O}(nd^{\prime})$ Proxy fidelity (task-dependent) Low
BADGE $\mathcal{O}(nd)$ High-$d$ gradient clustering High
SubSelNet $\mathcal{O}(n^{2}d)$ + meta-learning Quadratic in $n$ + training overhead High
[…]
Table 1: Comparison of subset selection methods. $n$: dataset size, $d$: feature dimension, $K$: batch size, $R$: selected rank/size, $d^{\prime}$: proxy model dimension.
Preliminaries
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a training dataset partitioned into $B$ mini-batches $\{\mathcal{X}_{i}\}_{i=1}^{B}$ of size $K\ll n$. For a batch $\mathcal{X}_{i}$, the gradient matrix $G\in\mathbb{R}^{d\times K}$ is defined as $G=[\nabla_{\Theta}L(\Theta;x_{1}),\dots,\nabla_{\Theta}L(\Theta;x_{K})]$. We project $\mathcal{X}_{i}$ into a low-rank embedding $V=f(\mathcal{X}_{i})\in\mathbb{R}^{K\times R}$ ($R\ll d$) and use Fast Maxvol to select $R$ rows $\mathbf{p}=[p_{1},\dots,p_{R}]$ maximizing submatrix volume. The subset $G_{R}=G(\mathrel{\mathop{\ordinarycolon}},\mathbf{p})$ approximates the full-batch gradient $\bar{g}$ via $\tilde{g}=\frac{1}{R}\sum_{r=1}^{R}G(\mathrel{\mathop{\ordinarycolon}},p_{r})$, with $\|\bar{g}-\tilde{g}\|_{2}\leq\varepsilon$.
3 Data Point selection via Row Sampling
The proposed methodology capitalizes on data point sampling by organizing the training dataset into a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, where each row $\mathbf{x}_{i}$ (for $i=1,\dots,n$) represents a distinct data point in the $d$-dimensional feature space. Relevant data points are identified by selecting a subset of rows from this matrix, a process facilitated by the Nyström method or analogous row/column subset selection techniques [13]. In contrast to traditional low-rank matrix approximation approaches, which focus on global matrix factorization, the column/row sampling strategy specifically seeks to identify a submatrix that preserves the critical structural information and intrinsic relationships embedded within the original matrix. A prevalent technique in this domain is the Cross-2D method [14], which leverages the Maxvol algorithm [15] to iteratively select both rows and columns in a manner that maximizes the volume of the submatrix formed at each step. Let $\mathcal{S}_{r}\subseteq\{1,\dots,n\}$ and $\mathcal{S}_{c}\subseteq\{1,\dots,d\}$ denote the set of row and column indices selected at any given iteration, respectively. The procedure initiates by selecting columns $\mathbf{x}_{j}$ that maximize the volume of the submatrix $\mathbf{X}[\mathcal{S}_{r},\mathcal{S}_{c}]$ formed by previously selected rows. Subsequently, rows are chosen iteratively based on the submatrix formed by the selected columns, thus refining the quality of the approximation in each iteration. This iterative process significantly enhances the matrix approximation by progressively selecting rows and columns that optimally capture the underlying structure of the data, minimizing the error in the low-rank approximation and ensuring computational efficiency. The result is a reduced-dimensional representation that maintains the essential properties of the original dataset while mitigating the computational burden associated with handling large-scale data matrices. Although the Cross-2D method is a well-established technique, the concurrent selection of both rows and columns introduces significant complexity into the selection process. Moreover, its performance is highly sensitive to the initial choice of row or column indices, with different initializations potentially yielding divergent results. Our experimental results, as detailed in Section 4 and illustrated in Figure 4, demonstrate that this method did not perform favorably in the context of our study.
3.1 Sampling of Ordered Extracted Features
Our approach simplifies the process by focusing solely on row sampling. Rather than sampling directly from the original matrix, we conduct accelerated sampling on a feature matrix, where the columns correspond to features ordered in descending importance. This methodology unfolds in two distinct stages, feature extraction and sample selection.
Step 1: Feature Extraction Let $A\in\mathbb{R}^{K\times M}$ denote a batch consisting of $K$ training samples, where each sample lies in an $M$-dimensional input space. A feature extraction function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{K\times M}\to\mathbb{R}^{K\times R}$ is applied to obtain a compact feature representation, producing the matrix $V=f(A)$, where $V\in\mathbb{R}^{K\times R}$ and $R\ll M$ is the number of retained features. Each row $\mathbf{v}_{k}=V_{k,\mathrel{\mathop{\ordinarycolon}}}\in\mathbb{R}^{R}$ represents the feature vector associated with the $k$-th sample. The columns of $V$ are ordered such that the features appear in descending order of relevance, with the most informative dimensions occupying the leftmost columns. Let $\text{Rel}(j)$ denote a feature relevance score for column $j$, computed using metrics such as variance, mutual information with the label space, or correlation. The ordering satisfies $\text{Rel}(1)\geq\text{Rel}(2)\geq\dots\geq\text{Rel}(R).$ The feature extraction function $f$ can be instantiated via several methods. One approach involves applying singular value decomposition (SVD) to $A$, yielding $A=U\Sigma V^{\top}$, and selecting the top $R$ left singular vectors $U_{R}\in\mathbb{R}^{K\times R}$ to form the feature matrix. Alternatively, $f$ may correspond to a nonlinear mapping $\phi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{M}\to\mathbb{R}^{R}$, such as the encoder of a trained autoencoder or a shallow neural network. Regardless of the specific implementation, the goal is to ensure that $V$ captures the dominant structural components of the data in a low-dimensional space suitable for efficient sample selection.
Step 2: Sample Selection using Fast MaxVol Method Given the feature matrix $\mathbf{V}\in\mathbb{R}^{K\times R}$, the goal is to sequentially select sample indices $p_{r}$ from the top-$R$ features such that the submatrix formed by the selected samples maximizes the volume, capturing the most significant information. One can apply the “conventional” Maxvol algorithm [15], which iterates the row index selection until the entries of the interpolation matrix are below a predefined threshold (often close to 1). Alternatively, we show a fast Max-volume algorithm with $R$ iterations to select $R$ row indices. Let $\mathbf{p}=[p_{1},\dots,p_{R}]$ denote the set of selected row indices.
  • Select the first data point index: Start with the first column $\mathbf{v}_{1}=\mathbf{V}(\mathrel{\mathop{\ordinarycolon}},1)$ and select the index $p_{1}$ corresponding to the element with the maximum absolute value, i.e., the most dominant entry in the most relevant feature $p_{1}=\arg\max_{i}\,|{\bf v}_{1}(i)|\,$
  • Selecting subsequent data point indices: The second data point index, $p_{2}$, is selected such that the submatrix ${\bf V}([p_{1},p_{2}],[1,2])$ has the maximum volume, i.e., maximum absolute value of the determinant $p_{2}=\arg\max_{i}|\text{det}({\bf V}([p_{1},i],[1,2]))|$ where the determinant of the submatrix ${\bf V}([p_{1},i],[1,2])$ is computed as
    $\text{det}({\bf V}([p_{1},i],[1,2]))$ $={\bf v}_{p_{1},1}{\bf v}_{i,2}-{\bf v}_{p_{1},2}{\bf v}_{i,1}$
    $={\bf v}_{p_{1},1}{\bf r}_{i,1}$ (1)
    Here the residual ${\bf r}_{1}={\bf v}_{2}-{\bf v}_{1}\left({\bf v}_{p_{1},1}\right)^{-1}{\bf v}_{p_{1},2}$. The above projection nullifies the $p_{1}$ element in ${\bf r}_{1}$, i.e., ${\bf r}_{1}(p_{1})=0$. From (1), finding the index $i$, which maximizes the volume of ${\bf V}([p_{1},i],[1,2]))$ in (3.1) is equivalent to identifying the maximum absolute element of ${\bf r}_{1}$, i.e., $p_{2}=\arg\max_{i}\,|{\bf r}_{1}(i)|$. We next update the selected index set ${\bf p}=[p_{1},p_{2}]$, and proceed with finding the other indices.
    Assume the selected index set ${\bf p}=[p_{1},p_{2},\ldots,p_{j-1}]$, we need to select the $j$-th column index such that the submatrix ${\bf V}([{\bf p},p_{j}],[1\mathrel{\mathop{\ordinarycolon}}j])$ of $j$ rows and the first $j$ columns has the maximum volume, i.e.,
    $p_{j}=\arg\max_{i}\,\left|\det\big{(}\mathbf{V}([\mathbf{p},i],[1\mathrel{\mathop{\ordinarycolon}}j])\big{)}\right|.$
    Suppose that ${\bf V}({\bf p},[1\mathrel{\mathop{\ordinarycolon}}j-1])$ is invertible, we define the residual term ${\bf r}_{j}={\bf v}_{j}-{\bf V}(\mathrel{\mathop{\ordinarycolon}},[1\mathrel{\mathop{\ordinarycolon}}j-1])\left({\bf V}({\bf p},[1\mathrel{\mathop{\ordinarycolon}}j-1])\right)^{-1}{\bf v}_{{\bf p},j}$. By applying Sylvester’s determinant theorem to the block matrix
    $\mathbf{V}([\mathbf{p},i],[1\mathrel{\mathop{\ordinarycolon}}j])=\begin{bmatrix}\mathbf{V}(\mathbf{p},[1\mathrel{\mathop{\ordinarycolon}}j-1])&\mathbf{v}_{\mathbf{p},j}\\ \mathbf{v}_{i,1\mathrel{\mathop{\ordinarycolon}}j-1}&\mathbf{v}_{i,j}\end{bmatrix}$,
    the determinant simplifies to $\det\big{(}\mathbf{V}([\mathbf{p},i],[1\mathrel{\mathop{\ordinarycolon}}j])\big{)}=\det\big{(}\mathbf{V}(\mathbf{p},[1\mathrel{\mathop{\ordinarycolon}}j-1])\big{)}\cdot\mathbf{r}_{j}(i).$ Hence, finding the most relevant index $p_{j}$ reduces to selecting the index corresponding to the maximum absolute value of ${\bf r}_{j}$ $p_{j}=\arg\max_{i}\,|\mathbf{r}_{j}(i)|.$
Remark 1 (Gradient Approximation Guarantee).
Let $A\in\mathbb{R}^{K\times M}$ be a batch matrix, and let $V=U_{R}\in\mathbb{R}^{K\times R}$ denote the matrix of the top-$R$ left singular vectors of $A$. Let $S\subset[K]$, $|S|=R$, be the subset of rows selected by the MaxVol algorithm applied to $V$. Since $V$ captures the dominant subspace of $A$, the MaxVol-selected rows in $V$ correspond to influential samples in $A$, ensuring their gradients approximate the full-batch gradient.If the gradient map $g(x)=\nabla_{\theta}L(\theta;x)$ is $L_{g}$-Lipschitz continuous, then: $\left\|\nabla_{\theta}L(\theta;A)-\nabla_{\theta}L(\theta;A(S,\cdot))\right\|_{2}\leq\frac{K}{R}L_{g}\sigma_{R+1},$ where $\sigma_{R+1}$ is the $(R+1)$-th singular value of $A$.
3.2 Dynamic Gradient based refinement
Input: Training dataset $\mathcal{X}$, ranks $\mathbf{R}=\{R_{i}\}_{i=1}^{c}$, feature matrices $\mathbf{V}$, selection period $S$, and batch size $K$
Output: Sampled data subset indices $\mathcal{S}^{t}$
for $t=1,\ldots,T$ do
 Stage 1: Subset Selection
 if $t\bmod S==0$ then
    for each batch $\mathcal{X}_{i}\subset\mathcal{X}$ do
       $\bar{\mathbf{g}}_{i}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\nabla_{\Theta}L(\Theta^{t};\mathcal{X}_{i}(\mathrel{\mathop{\ordinarycolon}},k))$
       for $R_{i}\in\mathbf{R}$ do
          $\mathcal{S}_{i}^{r}=\text{fast-maxvol}(\mathbf{V}_{i},R_{i})$
          $\mathbf{G}_{R_{i}}=\left[\ldots,\nabla_{\Theta}L(\Theta^{t};\mathcal{X}_{i}(\mathrel{\mathop{\ordinarycolon}},\mathcal{S}_{i}^{r}(p))),\ldots\right]$
          $d_{R_{i}}=\|\bar{\mathbf{g}}_{i}-\mathbf{G}_{R_{i}}\mathbf{G}_{R_{i}}^{\dagger}\bar{\mathbf{g}}_{i}\|_{2}^{2}$
       $R^{*}=\arg\min_{R_{i}}\{d_{R_{i}}\}_{i=1}^{|\mathbf{R}|}$
       $\mathcal{S}_{i}=\mathcal{S}_{i}^{R^{*}}$
    $\mathcal{S}^{t}=\{\mathcal{S}_{i}\}_{i=1}^{I}$
 else
    $\mathcal{S}^{t}=\mathcal{S}^{t-1}$
 Stage 2: Model Update
  Update all model parameters using $\mathcal{S}^{t}$
Algorithm 1 Training with Gradient-Aligned Sampling
A key challenge in the proposed algorithm is determining the optimal number of representative samples, or the number of top-$R$ features to select. This number, denoted as the input parameter $R$, must balance the size of the subset against the potential loss in accuracy. We propose selecting $R$ such that the gradient direction computed from the selected samples closely aligns with the gradient direction of the per-iteration full batch gradients. The rank $R$ can either be fixed or dynamically updated every $S$ training iterations, where $S$ is typically set to values between $20$ to $50$. In fixed-rank/fixed sample-size schemes, the subset size $R$ is predetermined using heuristics or a pre-defined metric such as in [5] and remains constant throughout training. However, this approach fails to consider the evolving nature of the optimization landscape. GRAFT dynamically selects $R$ from a set of candidates to minimize the projection error $\|\bar{g}-P_{R}\bar{g}\|_{2}^{2}$, selecting $R$ that satisfies a predefined error threshold $\epsilon$ [16].
Gradient Approximation with Sampled Subsets
The training dataset $X$ is randomly partitioned into batches $X_{i}$, each of size $m\times K$, where $K$ is the number of data points in the batch. The batch gradient $\bar{g}$ is calculated as the average gradient vector for the batch
$\bar{g}=\nabla_{\Theta}\mathcal{L}(\Theta;X_{i})=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\Theta}\mathcal{L}(\Theta;X_{i}(\mathrel{\mathop{\ordinarycolon}},k)).$
This gradient serves as the reference direction. Since gradients from fewer samples may result in less precise approximations, the goal is to select a minimal number of dominant samples $R$ such that their gradients, denoted $G_{R}=[g_{1},g_{2},\ldots,g_{R}]$, approximate the batch gradient. Here, $g_{r}=\nabla_{\Theta}\mathcal{L}(\Theta;X_{i}(\mathrel{\mathop{\ordinarycolon}},p_{r}))$ for $r=1,\ldots,R$. The key requirement is that $\bar{g}$ lies in the subspace spanned by $G_{R}$ $R=\arg\min_{R}d(\bar{g},G_{R}),$ where $d(\bar{g},G_{R})$ can be defined as angular error $d(\bar{g},G_{R})=\arcsin(\|\tilde{g}-\tilde{G}_{R}\tilde{G}_{R}^{T}\tilde{g}\|_{2})=\arcsin\left(\sqrt{1-\|\tilde{G}_{R}^{T}\tilde{g}\|_{2}^{2}}\right),$ or projection error $d(\bar{g},G_{R})=\|\bar{g}-\tilde{G}_{R}\tilde{G}_{R}^{T}\tilde{g}\|_{2}^{2}=1-\|\tilde{G}_{R}^{T}\tilde{g}\|_{2}^{2},$ Here, $\tilde{g}$ is the unit vector of the reference gradient $\bar{g}$, and $\tilde{G}_{R}$ is an orthogonal basis of the gradient matrix $G_{R}$ [16, 17].
Dynamic Rank Adjustment We propose dynamically selecting $R$ to control projection error $\|\bar{g}-P_{R}\bar{g}\|_{2}^{2}$. Specifically, we search over candidate ranks $R_{1}<\cdots<R_{m}$ and choose $R^{*}=\arg\min_{R_{i}}\|\bar{g}-P_{R_{i}}\bar{g}\|_{2}^{2},\quad\text{subject to}\quad\|\bar{g}-P_{R_{i}}\bar{g}\|_{2}^{2}\leq\epsilon,$ per-iteration batch during the selection iteration $S$. By Corollary 2, bounding the projection error ensures convergence under standard smoothness and variance assumptions [16].
Theorem 1 (Convergence via Gradient-Aligned Subspace Sampling).
Let $\bar{g}\in\mathbb{R}^{d}$ be the average gradient over a batch of $K$ samples, and $G_{R}\in\mathbb{R}^{d\times R}$ a subset of $R$ gradients selected via Maxvol, spanning a full-rank subspace $\mathcal{S}_{R}$. If the projection error $\|\bar{g}-\text{Proj}_{\mathcal{S}_{R}}(\bar{g})\|\leq\epsilon$, then gradient descent using the projected direction converges to a local minimum under standard assumptions [16].
Lemma 1 (Projection Error Bound).
Let $\tilde{G}_{R}$ be an orthonormal basis of $G_{R}$. Then
$\|\bar{g}-\tilde{G}_{R}\tilde{G}_{R}^{\top}\bar{g}\|_{2}^{2}=\|\bar{g}\|_{2}^{2}\left(1-\left\|\frac{\tilde{G}_{R}^{\top}\bar{g}}{\|\bar{g}\|_{2}}\right\|_{2}^{2}\right).$
Proof.
Decomposing $\bar{g}=\tilde{G}_{R}\tilde{G}_{R}^{\top}\bar{g}+r$, where $r\perp\text{span}(G_{R})$, the result follows by the Pythagorean theorem and the orthonormality of $\tilde{G}_{R}$ [18]. ∎
Corollary 1 (Dynamic Rank Adjustment Ensures Convergence).
If the rank $R$ is adjusted to keep $\|\bar{g}-\tilde{G}_{R}\tilde{G}_{R}^{\top}\bar{g}\|_{2}^{2}\leq\epsilon$, then GRAFT ensures convergence to a local minimum.
Proof.
From Lemma 3, bounding the residual energy implies $\left\|\frac{\tilde{G}_{R}^{\top}\bar{g}}{\|\bar{g}\|_{2}}\right\|_{2}^{2}\geq 1-\frac{\epsilon}{\|\bar{g}\|_{2}^{2}}.$ Hence, the projected direction retains sufficient gradient information, guaranteeing convergence under standard smoothness assumptions [16]. Please refer to Appendix for full proofs. ∎
(a) Alignment heatmap. Batch-wise cosine similarities increase over training epochs. † marks epochs with sufficient alignment ($\cos\theta>0.5$).
(b) Epoch-level trend. When $\cos\theta>0.5$, lower ranks $R^{*}$ can be selected while still satisfying the projection error constraint $\|\bar{g}-\mathbb{P}_{R}(\bar{g})\|_{2}^{2}\leq\epsilon$.
(c) Class Distribution. Bar plot illustrating the distribution of samples per class across timesteps.
Figure 2: Gradient alignment, rank adaptation, and class distribution during training. (a) captures per-batch variation, (b) summarizes global trends, and (c) visualizes class sampling distribution.
Gradient Alignment and Rank Evaluation
The gradient alignment dynamics, illustrated in Figures 2(a) and 2(b), provide empirical validation for GRAFT’s dynamic rank selection mechanism. Figure 2(a) shows that the cosine similarity $\cos\theta_{B_{t},E_{t}}$ between batch-level and epoch-level gradients improves from 0.53 to 0.89 during training, with alignment values exceeding 0.5 (marked ’+’) in the majority of cases. This ensures that the projection error satisfies $\|\mathbf{r}\|_{2}^{2}<0.75\|\mathbf{g}\|_{2}^{2}$ (Lemma 3), thereby meeting the $\epsilon$-approximation guarantees described in Corollary 2. The global alignment statistics in Figure 2(b) further reveal stable behavior, with mean $\mu=0.72$ and standard deviation $\sigma=0.15$. Over 95% of alignment values fall within the interval [0.42, 1.02][0.42,\ 1.02], indicating that the selected subsets consistently preserve the dominant gradient direction. Notably, the strong correlation between high alignment and rank reduction confirms that GRAFT adaptively optimizes the subset size $R$ without compromising convergence. Rare instances of low alignment ($\cos\theta<0.5$) are effectively mitigated by the dynamic adjustment mechanism. This gradient-aligned selection process preserves the balanced class representation, Figure  2(c) shows how the method dynamically distributes samples across classes throughout training.
3.3 Complexity Analysis
We establish convergence guarantees for GRAFT under standard assumptions: (A1) $\mathcal{L}$ is $L$-smooth; (A2) stochastic gradients are unbiased with bounded variance; (A3) projection error satisfies $\|\bar{g}_{i}-G_{R}G_{R}^{\dagger}\bar{g}_{i}\|_{2}\leq\varepsilon$.
Theorem 2 (Projected-gradient convergence).
Under (A1)–(A3), SGD with projected directions $G_{R}G_{R}^{\dagger}\bar{g}_{i}$ converges to a stationary point with $\min_{t\leq T}\mathbb{E}\big{\|}\nabla\mathcal{L}(\Theta^{t})\big{\|}_{2}^{2}\leq O\!\left(\frac{1}{\sqrt{T}}\right)+O(\varepsilon^{2})$
Theorem 3 (Computational complexity).
Fast MaxVol selection with gradient-projection evaluation costs $O(Kr_{\max}d_{z})+O(KR^{2})+O(pR^{2})$, where $K$ is batch size, $R$ is selected rank, $d_{z}$ is feature dimension, and $p=|\Theta|$. Cost is independent of dataset size $N$.
4 Experiments
Figure 3: Performance-efficiency trade-offs showing (top) normalized accuracy $\Phi_{\text{acc}}(\mathcal{E})$ vs. CO2 emissions $\mathcal{E}$ and (bottom) data utilization $\Psi(f)$ vs. fraction $f$ for five datasets. Solid curves show fitted exponential gain functions $E_{0}+(H-E_{0})(1-e^{-\lambda x/x_{\max}})$. GRAFT variants achieve higher $\Phi_{\text{acc}}$ at lower $\mathcal{E}$ and maintain $\Psi(f)>0.8H$ with 60% less data than conventional methods.
We developed a rigorous framework to evaluate the fundamental trade-offs between model performance and computational efficiency across dataset reduction methods. The analysis characterizes two key relationships (1) the fidelity coefficient $\Phi_{\text{acc}}(\mathcal{E})={\text{Accuracy}(\mathcal{E})}/{\text{Accuracy}_{\text{(full)}}}$ as a function of CO2 emissions $\mathcal{E}$, and (2) the data utilization factor $\Psi(f)={\text{Accuracy}(f)}/{\text{Accuracy}_{\text{(full)}}}$ as a function of dataset fraction $f\in[0,1]$. The empirical results for the fidelity coefficient and data utilization factor are visualized in 3. The normalized metrics enable direct comparison across methods and datasets. We compared GRAFT with four integrated subset selection baselines and one preselection method (1) full data training, (2) CRAIG [7], (3) GLISTER [6], (4) GradMatch [5], and (5) DRoP [4]. The performance trajectories were modeled using a generalized exponential gain function $E(x)=E_{0}+(H-E_{0})(1-e^{-\lambda x/x_{\max}})$ where $E_{0}$ represents baseline performance at minimal resource investment, $H$ is the asymptotic maximum performance, $\lambda$ controls the convergence rate, and $x_{\max}$ provides normalization. Maximum likelihood estimation was performed separately for each method-dataset combination, with goodness-of-fit quantified through the coefficient of determination $R^{2}=1-{(\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}/{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2})}$ The CO2 emissions were estimated using the [19] library, which provides real-time tracking of energy consumption and its environmental impact. The calculations account for factors such as hardware specifications, duration of model training, and the carbon intensity of the local energy grid. Specifically, emissions are calculated as $\mathcal{E}=P\times t\times I,$ where $P$ denotes the power consumption of the hardware in kilowatts (kW), $t$ represents the duration of model training in hours (h), and $I$ is the carbon intensity of the energy grid in kgCO2/kWh.
Analysis GRAFT demonstrates superior performance characteristics across all metrics. For $\Phi_{\text{acc}}(\mathcal{E})$, GRAFT achieved $\lambda$ values 1.8–2.4× higher than competing methods (mean $\lambda=3.2\pm 0.4$ vs. $1.4\pm 0.3$ for GRADMATCH), indicating faster accuracy gains per unit CO2. The warm-start variant GRAFT Warm showed particularly strong performance at low emissions $(\mathcal{E}<0.05\,\text{kg})$, with $E_{0}=0.38\pm 0.05$ compared to $0.12\pm 0.08$ for standard GRAFT, demonstrating the benefit of initialization with full-data representations.The data utilization analysis revealed $\Psi(f)$ curves where GRAFT methods maintained $\Psi(f)>0.8$ at $f=0.25$, while other methods required $f\geq 0.4$ to reach equivalent performance levels. This 60% reduction in required training data stems from GRAFT’s dual optimization of both gradient matching and representational diversity. The $R^{2}$ values confirmed excellent model fits across all conditions (GRAFT: $0.92\pm 0.03$, GRAFT Warm: $0.94\pm 0.02$), validating the exponential gain model’s appropriateness for characterizing these relationships. Notably, the performance gap between GRAFT variants and other methods increased with decreasing $f$, with relative accuracy improvements following $\Delta\Psi(f)\approx 0.25{(1-f)}/{f}\quad\text{for}\quad f\in(0,0.5].$
This relationship suggests that GRAFT’s advantages become increasingly pronounced in low-data regimes, making it particularly suitable for resource-constrained applications. The consistent alignment of these empirical results provides strong evidence for GRAFT’s superior sample efficiency and computational optimality. Across all datasets, GRAFT methods consistently minimized CO2 emissions while achieving higher accuracy. For instance, in CIFAR10 [20], GRAFT achieved a $0.15\,\text{kg}$ CO2 reduction compared to GRADMATCH at a similar fidelity level. In TinyImagenet [21], the gap was even more pronounced, with GRAFT reducing emissions by $0.21\,\text{kg}$ on average for equivalent accuracy. In the case of Caltech256  [22], GRAFT demonstrated substantial gains, cutting CO2 emissions by $0.28\,\text{kg}$ compared to the nearest competitor while achieving $0.631$ fidelity with just $0.15$ fraction of the dataset.
Table 2: CO2 Emissions (kg) and Accuracy (%) for BERT on IMDB.
Method Emiss (Kg) Top-1 Acc (%)
Full (Baseline) 0.32 93.92
GRAFT (10%) 0.05 91.72
GRAFT Warm (10%) 0.14 93.74
GRAFT (35%) 0.15 93.56
GRAFT Warm (35%) 0.19 93.71
Fine-tuning transformers The results in  2 demonstrate GRAFT Warm’s accuracy superiority over standard GRAFT at equivalent data fractions during transformer fine-tuning. At 35% data, GRAFT Warm achieves 93.71% accuracy (vs. 93.56% for GRAFT) with only 0.04 kg higher emissions (0.19 kg vs. 0.15 kg). This 0.15% accuracy gain demonstrates the effectiveness of warm-start initialization, particularly notable given: (1) the 0.18% gap to full-dataset performance (93.71% vs. 93.92%) becomes negligible in practice, and (2) the 41% emission reduction (0.19 kg vs. 0.32 kg) remains substantial. The pattern holds at 10% data, where GRAFT Warm reaches 93.74% accuracy (vs. 91.72% for GRAFT) by leveraging pretrained representations, albeit with higher emissions (0.14 kg vs. 0.05 kg). This trade-off suggests warm-starting is preferable when maximal accuracy is critical, while cold-start GRAFT offers better efficiency for moderate accuracy targets particularly useful for transformer finetuning.
Figure 4: Ablation study on CIFAR-10: (Left) test accuracy for different feature extractors (SVD, AE, ICA). (Right) convergence of sample selection methods (FastMaxVol vs. CrossMaxVol). Shaded regions indicate $\pm$1 std. dev. across three runs.
Ablation We evaluated GRAFT on CIFAR-10 using a ResNet-18 backbone by comparing three feature extraction methods - Singular Value Decomposition (SVD), an autoencoder (AE), and Independent Component Analysis (ICA). SVD decomposed the batch matrix $A\in\mathbb{R}^{K\times M}$ into its top-$R$ singular vectors, where $K$ is the batch size and $M$ is the flattened image dimension, while the AE used a shallow encoder-decoder to learn latent representations, and ICA identified features based on variance contributions. Results in Figure 4 demonstrated that SVD outperformed both AE and ICA when using only 25% of the training data, achieving a test accuracy of 90.3%, compared to 86.7% for AE and 80.8% for ICA, reflecting SVD’s strength in preserving dominant data structures. To further explore efficiency and scalability, we assessed the three extractors with a lightweight logistic regression classifier on CIFAR-10, running three trials with seeds 42–46 and a fixed 20% holdout test set. As illustrated in Table 4, AE obtained the highest accuracy of 38.99% but incurred a 5× higher computational cost (0.1996 s/batch) compared to SVD, which balanced competitive accuracy (38.19%) with much faster processing (0.0385 s/batch), while ICA was the slowest (0.3921 s/batch) despite moderate accuracy (38.25%).
Table 3: Feature extraction performance (mean $\pm$ std over 5 runs)
Method Acc (%) Time (s/batch) Significance
SVD ($R=64$) $38.19\pm 0.24$ $0.0385\pm 0.0025$
AE $38.99\pm 0.21$ $0.1996\pm 0.0105$ $p=0.0066$
ICA $38.25\pm 0.0$ $0.3921\pm 0.0512$ $p=0.6483$
Table 4: Similarity & Speed
Method Similarity Time (s)
Fast MaxVol 0.6250 0.0005
CrossMaxVol 0.5938 0.0456
Subspace Similarity and Efficiency The results presented in Table 4 compare the Subspace Similarity and Execution Time for the proposed Fast MaxVol and the baseline CrossMaxVol, implemented via the teneva [23] library on the Iris dataset [24]. Subspace similarity is quantified as the sum of squared cosines of the principal angles between the subspaces spanned by the selected samples $(V_{1},V_{2})=\sum_{i=1}^{k}\cos^{2}(\theta_{i})$, where $V_{1}$ and $V_{2}$ are the subspace bases, and $\theta_{i}$ are the principal angles. Fast MaxVol achieves a similarity of $0.625$, outperforming CrossMaxVol ($0.59375$), indicating that the subspace spanned by Fast MaxVol is more aligned with the optimal representation. Additionally, Fast MaxVol also converges faster, as illustrated in 4 (right) on ResNet-18, where the method demonstrates faster convergence and execution time of $0.000539$ seconds compared to $0.045594$ seconds for CrossMaxVol in table  4, which represents a speedup of approximately $84.6\times$.
5 Limitations and Future Work
While GRAFT offers significant efficiency gains, its performance depends on the quality of low-rank feature extraction, and suboptimal feature representations may degrade gradient approximation. The dynamic rank adjustment mechanism, though effective, introduces computational overhead. Additionally, the method’s reliance on batch-level subspace approximations may limit its robustness in highly non-IID or imbalanced data settings. Future research could explore the application of Fast MaxVol beyond data subset selection, such as in channel pruning for neural networks. By selecting the most informative channels in each layer, Fast MaxVol could potentially reduce the computational cost of inference without significant accuracy loss. Table 5 shows a preliminary result of Fast MaxVol used to prune 50% of channels in a ResNet-18 model.
Table 5: Performance of Fast MaxVol for channel pruning (ResNet-18, CIFAR-10, Selective, 50%)
Method Parameters(Million) Accuracy (%) GFLOPs Inference Time(sec)
Baseline 11.18 93.21 1.82 6.74
Fast MaxVol 5.61 91.97 1.09 3.79
6 Conclusion
We present GRAFT, a gradient-aware data sampling framework that strategically leverages fast Max-Volume and gradient alignment to dynamically select representative samples during training. Unlike traditional pre-selection approaches, GRAFT is fully integrated into the training loop, ensuring that selected subsets both preserve maximal volume in the feature space and align closely with the global gradient trajectory. This dual-focus sampling accelerates convergence, reduces training times, and substantially optimizes energy consumption and CO2 emissions all with minimal accuracy loss. The proposed framework is particularly well-suited for resource-constrained training, hyperparameter optimization (HPO), and AutoML pipelines, offering a scalable solution to sustainable machine learning.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s contributions, including the proposed GRAFT framework for efficient data subset selection, its gradient-aware sampling mechanism, and its benefits in reducing computational cost and environmental impact. The claims are supported by theoretical analysis and experimental results across multiple datasets.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a dedicated "Limitations and Future Work" section (Section 5.1) that discusses several limitations, such as reliance on the quality of low-rank feature extraction, computational overhead from dynamic rank adjustment, and sensitivity to batch size and feature extraction methods.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides theoretical results (Theorems 1-4, Lemmas 1-3) with clear assumptions and complete proofs in the main text and Appendix A. The proofs are mathematically rigorous and well referenced.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides extensive experimental details, including datasets used (CIFAR-10, CIFAR-100, FashionMNIST, TinyImageNet, IMDB, DermanMNIST), model architectures (ResNet-18, ResNeXt29, MobileNetV2, BERT), hyperparameters (learning rates, batch sizes, weight decay), and evaluation metrics. The Appendix includes implementation details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The paper states that "The code implementation for GRAFT will be provided after acceptance," but no code or data is currently available in the submitted version.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper provides comprehensive experimental settings, including optimizer details (SGD with Nesterov momentum), learning rate schedules (CosineAnnealing), batch sizes, and data fractions used for subset selection. All details are in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports mean and standard deviation over multiple runs (e.g., Tables 2-3, Figure 5) and includes statistical significance tests (p-values) where applicable.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper specifies hardware (Tesla V100-SXM2 GPU, Intel Xeon Gold CPU), memory (16GB), and execution times for key experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms to ethical guidelines. The paper focuses on algorithmic improvements for efficient training without involving sensitive data or human subjects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper highlights positive impacts like reduced computational costs and environmental benefits. While not explicitly discussing negative impacts, the method’s general-purpose nature doesn’t suggest specific misuse risks beyond typical ML applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The research does not involve high-risk models or datasets that would require special safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets used (CIFAR, FashionMNIST, etc.) are properly cited, and standard academic licenses are assumed for these widely-used benchmarks.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new datasets or models are being released with this submission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The research does not involve human subjects or crowdsourcing.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The research does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used in developing the core methodology of this research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.